# A One Dimensional (1D) Computational Fluid Dynamics Study of Fontan-Associated Liver Disease (FALD)

### v1

@@ 2.2 One Dimensional Fluid Dynamics Model
Similarly to [46, 36, 49, 48], we predict pressure $p(x,t)$ (mmHg), blood flow $q(x,t)$ (mL/s), and area deformation $A(x,t)$ (cm2) in each vessel within the portal vein and hepatic artery vasculatures using a 1D fluid dynamics model. The model is derived from the Navier-Stokes equations, which assume that blood flow is axisymmetric and Newtonian. We also assume that blood is incompressible, viscous, and homogeneous with constant density $\rho$, and that each vessel can be represented by a cylinder with an impermeable wall. Under these assumptions, mass conservation and momentum balance are governed by
@@
We impose a flat velocity profile with a linearly decreasing boundary layer with thickness $\delta=\sqrt{\nu T/2\pi}$ (cm) [50, 51, 52, 49], where $T$ (s) is the length of the cardiac cycle.
@@
where $E$ (g/cm/s²) is Young’s modulus, $h$ (cm) is the vessel wall thickness, $p_{0}$ (mmHg) is the reference pressure, $r_{0}$ (cm) is the inlet radius, and $A_{0}$ (cm²) is the cross-sectional area when the pressure equals its reference value [53]. We assume that the vessels become stiffer as they get smaller (equation (4), where $k_{1}$ (g/cm/s2), $k_{2}$ (1/cm), and $k_{3}$ (g/cm/s2) are constants obtained from literature values [47, 54, 36].
@@
where $p$ is a parent vessel and $d_{1}$ and $d_{2}$ are daughter vessels. Similarly to previous studies [46, 36, 49, 48], we solve the model using the two-step Lax-Wendroff finite difference scheme using.
At the outlet of the terminal vessels, we use a structured tree boundary condition prescribing flow to the small vessels in the vascular bed [55, 49, 36, 56]. Equations in the structured tree are obtained semianalytically by solving a linearized wave equation ignoring inertial forces, which is appropriate since viscous forces are dominant in small vessels.
The resulting wave equations stated in the frequency domain, described in detail in [49], are given by
@@
where $c=f(F_{J},C)$ denotes the wave propagation velocity [49]. The function $F_{J}(J_{0},J_{1})$ is a function of $J_{0}$ and $J_{1}$, the zeroth- and first-order Bessel functions, and $C$ is the vessel compliance. Solving the equation for flow $Q$ and pressure $P$, we get
@@
In this equation $a$ and $b$ are integration constants [36, 47, 49, 48].
@@
where $Z_{p},Z_{d_{1}}$, and $Z_{d_{2}}$ are the impedance values of the parent and daughter vessels. At the end of the structured tree, we assume that the terminal impedance at the end of the structured tree is zero. The equations in this tree are solved recursively for each vessel, predicting the impedance at the beginning of the vessel as a function of the impedance at the root of the vessel. The impedance at the root of the structured tree is transformed to the time domain by convolution. Following Riemann invariants, the root impedance forms the outflow boundary condition for large vessels.
@@ 2.3 Wall Shear Stress Approximation
$\tau_{w}=\left\{\begin{array}[]{lr}0,&r<R-\delta\\ \frac{\mu\bar{u}}{\delta},&R-\end{array}\right.$ (12)
where $\mu$ is blood viscosity, and $\delta=\sqrt{\frac{\nu T}{2\pi}}$ is the boundary layer thickness [36, 48].
Table 2: Patient-specific parameter values. PV refers to the portal vein, HA to the hepatic artery.
@@
Table 3: Global parameter values for all patients.
@@ 2.4 Perfusion
The perfusion and pressure distribution is determined by mapping the 1D flow and pressure predictions from the vessels’ terminals to the liver volume [57]. The end of each terminal vessel is denoted by the $x,y,z$ coordinate extracted from the centerlines. Each of these points is assigned the mean flow and pressure from the vessel in question. The flow and pressure are distributed to the volume using Matlab’s nearest-neighbor algorithm, providing a 3D representation of blood perfusion and pressure distribution in the liver. To accurately compare differences between patients with different inflows, the flow at the terminals of the hepatic and portal networks are normalized by the average inlet flow of the DORV patient.
@@ 2.5 Simulations
We simulate liver hemodynamics for the DORV and HLHS networks. To mimic FALD in the HLHS patient, we increase the vascular stiffness in both large and small vessels. For each network, the inflow propagates downstream through the large vessels of the respective portal vein and hepatic artery systems. We define four patient types,
  • a control DORV patient,
  • an HLHS patient labeled “healthy HLHS”,
  • an HLHS patient with early FALD progression, and
@@
To predict the liver hemodynamics in each patient, we calibrate the models using the dimensions in Table 2 and Table 3. We manually tune the nominal stiffness parameters of the large vessel for each patient to generate flow waveforms and pressure predictions that match both the data of the imaged region and previous studies [49, 58].
In this study, the stiffness of the small vessels, ${k}_{\text{sn}}$, is set equal to their counterparts of the large vessels, ${k}_{\text{n}}$ [36, 49].
@@ 2.5.2 FALD Simulation
Patients with liver cirrhosis experience increased vessel stiffness due to alterations in the composition of the extracellular matrix (ECM) [59, 60]. To predict the change in hemodynamics in the diseased patient under the development of FALD, we simulate a liver diseased with FALD by increasing downstream vascular resistance. We increase the parameter $fs_{3}$, resulting in a stiffness of 3.0 x 10 6(g/cm/s2) in the portal vein and 1.0 x 10 6(g/cm/s2) in the hepatic artery. Furthermore, we decrease the outlet boundary condition $r_{m}in$ to $0.005$ cm to further increase the resistance.
@@ 3 Results
We compare the flow and pressure predictions of a DORV, HLHS, and simulated HLHS patient with FALD to investigate differences in hemodynamics. Our results show that the HLHS patient has higher portal vein pressures than the DORV patient and that pressure reaches hypertensive levels under FALD conditions. We see that WSS is higher in the HLHS patient and increases with FALD. The perfusion plots show that the DORV patient has more evenly distributed flow throughout the tissue. It also shows a change in the flow distribution as FALD conditions increase.
@@ 3.1 Pressure and flow
Figure 3: Pressure and flow predictions for the DORV, HLHS, and simulated FALD HLHS patient in the portal network. The patient with simulated FALD is hypertensive and displays more reflective waves with increased magnitude.
Portal vein. Table 4 shows the predicted pressure and flow in the portal vein network. The average pressure is calculated as the mean pressure during the cardiac cycle. The HLHS patient has a higher systolic pressure, pulse pressure, and average flow than the DORV patient. In patients with simulated FALD HLHS, systolic pressures and pulse pressures increase significantly. The average flow remains the same as we did not alter the amount of blood entering each network.
Table 4: Portal vein pressure and flow predictions for the HLHS and DORV patients. The Pulse pressure denotes the range between systolic and diastolic pressures.
Healthy HLHS Early FALD Late FALD DORV
Systolic Pressure (mmHg) 17.91 20.19 22.9 16.29
Diastolic Pressure (mmHg) 4.76 3.72 2.24 4.25
Pulse (mmHg) 13.14 16.46 20.68 12.04
Average Pressure (mmHg) 10.73 10.90 10.93 10.20
Average Flow (mL/s) 15.85 15.85 15.85 11.85
@@
The pressure and flow of the inlet vessel (vessel 0) differ significantly between the two patients. The DORV patient has smaller reflective waves than the HLHS patient, and the magnitude and frequency of the reflective waves increase with the severity of FALD. The incident pressure wave becomes steeper and the increase in wave speed causes the reflected wave to coincide with the systolic peak; as a result, the pressure wave loses the dicrotic notch as FALD progresses.
@@
Figure 4: Pressure and flow predictions for vessels 0 and 13 in the hepatic artery vasculature in a DORV, HLHS, and FALD HLHS patient. FALD conditions shift the HLHS patient’s peak flow to earlier in the cardiac cycle in vessel 13 and raise the maximum flow slightly.
A similar analysis is performed for the hepatic artery network. Table 5 shows the predicted pressure and flow in the hepatic artery network. The HLHS patient has a higher systolic and pulse pressure compared to the DORV patient. However, there is no significant change in pressures under FALD conditions for the HLHS patient.
We selected two representative vessels (0, 13) in the vasculature of the liver artery to compare flow and pressure predictions between patients 4. Vessel 0 is the hepatic artery, the network’s root vessel, while vessel 13 is a terminal vessel. The pressure curves are steeper in the HLHS patient, with no significant changes as FALD is imposed. The flow curves change slightly under FALD conditions, with late-stage FALD shifting the peak of the curve to a somewhat earlier time in the cardiac cycle and presenting a higher maximum flow.
Table 5: Predicted pressure and flow in the hepatic artery network of the HLHS and DORV patients. The Pulse pressure denotes the range between systolic and diastolic pressures.
Healthy HLHS Early FALD Late FALD DORV
Systolic Pressure (mmHg) 100.14 99.87 100.02 97.30
Diastolic Pressure (mmHg) 74.64 74.28 74.33 77.56
Pulse (mmHg) 25.50 25.59 25.69 19.73
Average Pressure (mmHg) 86.36 86.30 86.25 85.62
Average Flow (mL/s) 6.79 6.79 6.79 5.08
@@ 3.2 Wall Shear Stress Approximation
Graphical representations of WSS over one cardiac cycle at the midpoint of the six previously selected vessels are shown in Figure 5. Inlet vessels (vessel 0) of both vasculatures show the most significant difference in WSS between patients. The HLHS patient not only has higher mean WSS values, but the maximum WSS within all vessels is greater than the DORV patient. These values increase with the degree of FALD. The peak of the WSS curve also shifts earlier in the cycle with increasing FALD progression.
Table 6: Average WSS in select DORV and HLHS patient vessels under normal and FALD conditions.
@@
Hepatic Artery (0) 48.16 47.92 48.17 34.53
Hepatic Artery (13) 22.12 22.04 21.90 15.72
Portal Vein (0) 5.95 5.91 5.90 4.27
Portal Vein (54) 1.02 0.98 0.98 0.74
Portal Vein (58) 1.45 1.31 1.32 1.04
Portal Vein (69) 0.40 0.33 0.34 0.29

### v2

@@ 2.2 One Dimensional Fluid Dynamics Model
Similarly to [Colebank19, alyssa2022, olufsen2000numerical, bartolo2023framework], we predict pressure $p(x,t)$ (mmHg), blood flow $q(x,t)$ (mL/s), and area deformation $A(x,t)$ (cm2) in each vessel within the portal vein and hepatic artery vasculatures using a 1D fluid dynamics model. The model is derived from the Navier-Stokes equations, which assume that blood flow is axisymmetric and Newtonian. We also assume that blood is incompressible, viscous, and homogeneous with constant density $\rho$, and that each vessel can be represented by a cylinder with an impermeable wall. Under these assumptions, mass conservation and momentum balance are governed by
@@
In the organ network studied here, the vessels are relatively short compared to their length. As a result, we do not anticipate a fully developed flow. To account for this, we impose a Stokes velocity profile with a linearly decreasing boundary layer with thickness $\delta=\sqrt{\nu T/2\pi}$ (cm) [nichols1991mcdonald, pedersen1993two, pries1992blood, olufsen2000numerical], where $T$ (s) is the length of the cardiac cycle.
@@
where $E$ (g/cm/s²) is Young’s modulus, $h$ (cm) is the vessel wall thickness, $p_{0}$ (mmHg) is the reference pressure, $r_{0}$ (cm) is the inlet radius, and $A_{0}$ (cm²) is the cross-sectional area when the pressure equals its reference value [qureshi2019hemodynamic]. We assume that the vessels become stiffer as they get smaller (equation (4), where $k_{1}$ (g/cm/s2), $k_{2}$ (1/cm), and $k_{3}$ (g/cm/s2) are constants obtained from literature values [Colebank21, paun2020assessing, alyssa2022].
@@
where $p$ is a parent vessel and $d_{1}$ and $d_{2}$ are daughter vessels. Similarly to previous studies [Colebank19, alyssa2022, olufsen2000numerical, bartolo2023framework], we solve the model using the two-step Lax-Wendroff finite difference scheme.
At the outlet of the terminal vessels, we use a structured tree boundary condition prescribing flow to the small vessels in the vascular bed [colebank2021sensitivity, olufsen2000numerical, alyssa2022, chambers2020structural]. Equations in the structured tree are obtained semianalytically by solving a linearized wave equation ignoring inertial forces, which is appropriate since viscous forces are dominant in small vessels.
The resulting wave equations stated in the frequency domain, described in detail in [olufsen2000numerical], are given by
@@
where $c=f(F_{J},C)$ denotes the wave propagation velocity [olufsen2000numerical]. The function $F_{J}(J_{0},J_{1})$ is a function of $J_{0}$ and $J_{1}$, the zeroth- and first-order Bessel functions, and $C$ is the vessel compliance. Solving the equation for flow $Q$ and pressure $P$, we get
@@
In this equation $a$ and $b$ are integration constants [alyssa2022, Colebank21, olufsen2000numerical, bartolo2023framework].
@@
where $Z_{p},Z_{d_{1}}$, and $Z_{d_{2}}$ are the impedance values of the parent and daughter vessels. It is assumed that the terminal impedance at the end of the structured tree is zero. The equations in this tree are solved recursively for each vessel, predicting the impedance at the beginning of the vessel as a function of the impedance at the root of the vessel. The impedance at the root of the structured tree is transformed to the time domain by convolution. Following Riemann invariants, the root impedance forms the outflow boundary condition for large vessels.
@@ 2.3 Wall Shear Stress Approximation
[…]
where $\mu$ is blood viscosity, and $\delta=\sqrt{\frac{\nu T}{2\pi}}$ is the boundary layer thickness [alyssa2022, bartolo2023framework].
Table 2: Patient-specific parameter values. PV refers to the portal vein, HA to the hepatic artery. $k_{1,2,3}$ denote large vessel stiffness, $ks_{1,2,3}$ denote small vessel stiffness, $p_{0}$ is the diastolic pressure coefficient, $T$ is length of one cardiac cycle, and $r_{\text{min}}$ denotes the terminal radius value that the structured tree is propagated to.
@@
Table 3: Global parameter values for all patients. $\rho$ denotes density, $\mu$ is viscosity, $\alpha$ and $\beta$ are structured tree parameters determining the asymmetry of the tree, and lrrlrr is the length-to-radius ratio of the structured tree.
@@ 2.4 Perfusion
The perfusion and pressure distribution is determined by mapping the 1D flow and pressure predictions from the vessels’ terminals to the liver volume [Colebank_perf]. The end of each terminal vessel is denoted by the $x,y,z$ coordinate extracted from the centerlines. Each of these points is assigned the mean flow and pressure from the vessel in question. The flow and pressure are distributed to the volume using Matlab’s nearest-neighbor algorithm, providing a 3D representation of blood perfusion and pressure distribution in the liver. To accurately compare differences between patients with different inflows, the flow at the terminals of the hepatic and portal networks are normalized by the average inlet flow of the DORV patient.
@@ 2.5 Simulations
Figure 4: Workflow chart. The number (1-8) corresponds to simulations described in 2.5. DORVn refers to the DORV patient network geometry, HLHSn refers to the HLHS patient network geomtery, and DORVp refers to the DORV parameters listed in Table 2. Figures of results for simulations 5-8 can be found in the supplement.
We simulate liver hemodynamics for the DORV and HLHS networks. To mimic FALD in the HLHS patient, we increase the vascular stiffness in the large and small vessels. For each network, the inflow propagates downstream through the large vessels of the portal vein and hepatic artery networks. We define four patient types:
  1. 1.
    A control DORV patient based on patient data.
  2. 2.
    A HLHS patient based on patient data.
  3. 3.
    An HLHS patient with early FALD progression.
  4. 4.
@@
To study the impact of inflow waveform, network geometry, and model parameters, we also include the following simulations (results shown in supplement):
  1. 5.
    HLHS inflow waveform effects: We imposed the HLHS inflow on the DORV and the HLHS networks, respectively, keeping the vascular parameters and patient specific geometry fixed,
  2. 6.
    DORV inflow waveform effects: We imposed the DORV inflow on the DORV and the HLHS networks, respectively, keeping the vascular parameters and patient specific geometry fixed,
  3. 7.
    Parameter effects: Fixing the inflow and network geometry we imposed the DORV parameters on the DORV and HLHS networks.
  4. 8.
    Portal vein Inflow waveform amplitude: As reported by [chavan], the inflow waveform to the portal vein is often blunted. To study effects of this waveform shape, we imposed digitized the waveform from [chavan] scaling it to the means for the two patients in this study. These simulations were conducted with the patient parameters reported in Table 2.
Results for simulations 5-8 are reported in the supplemental material.
@@
To predict the liver hemodynamics in each patient, we calibrate the models using the dimensions in Table 2 and Table 3. We manually tune the nominal stiffness parameters of the large vessel for each patient to generate flow waveforms and pressure predictions that give physically meaningful results, with average pressures and flows being comparable to literature values [olufsen2000numerical, garcia-pagan1996physical].
In this study, the stiffness of the small vessels, ${k}_{\text{sn}}$, is set equal to their counterparts of the large vessels, ${k}_{\text{n}}$ [alyssa2022, olufsen2000numerical].
@@ 2.5.2 FALD Simulation
Patients with liver cirrhosis experience increased vessel stiffness due to alterations in the composition of the extracellular matrix (ECM) [arriazu2014extracellular, villela2016NAFLD]. To predict the change in hemodynamics in the diseased patient under the development of FALD, we simulate a liver diseased with FALD by increasing downstream vascular resistance. We increase the parameter $fs_{3}$, resulting in a stiffness of 3.0 x 10 6(g/cm/s2) in the portal vein and 1.0 x 10 6(g/cm/s2) in the hepatic artery. Furthermore, we decrease the outlet boundary condition $r_{m}in$ to $0.005$ cm to further increase the resistance.
@@ 3 Results
We compare the flow and pressure predictions of a DORV, HLHS, and simulated HLHS patient with FALD to investigate differences in hemodynamics. We also investigate the impact of cardiac output and inflow waveform by imposing patient inflows on their counterpart network. Our results show that the HLHS patient has higher portal vein pressures than the DORV patient and that pressure reaches hypertensive levels under FALD conditions. These same trends are seen regardless of the inflow waveform imposed. We see that WSS is higher in the HLHS patient and increases with FALD. The perfusion plots show that the DORV patient has more evenly distributed flow throughout the tissue. It also shows a change in the flow distribution as FALD conditions increase.
@@ 3.1 Pressure and flow
Figure 5: Pressure and flow predictions for the DORV, HLHS, and simulated FALD HLHS patient in the portal network. We see that as the degree of FALD increase, the peak of the flow waveform is slightly shifted to the left, with systolic pressures slightly increasing.
3.1.1 Patient-specific inflows
Portal vein. Table 4 shows the predicted pressure and flow in the portal vein network. The average pressure is calculated as the mean pressure during the cardiac cycle. All patient types have approximately the same systolic pressure, with the DORV patient having a lower diastolic pressure. Naturally, the DORV patient also has the highest pulse pressure of the patient groups. However, of all patient types, the DORV patient has the lowest average pressure. In patients with simulated FALD HLHS, systolic pressures and pulse pressures increase slightly. The Late FALD patient has the highest systolic pressure and pulse pressure of the three HLHS patients. Regarding flow, FALD conditions shift the peak of the flow waveform to a somewhat earlier time in the cardiac cycle in the downstream vessels (Figure 5).
Table 4: Portal vein pressure predictions for the HLHS and DORV patients. All pressures shown have units mmHg. The Pulse pressure refer to the range between systolic and diastolic pressures. All values are listed in mmHg. Note, Sim 5 refers to the HLHS inflow imposed on DORV network, Sim 6 to the DORV inflow imposed on HLHS network, Sim 7 to the DORV parameters imposed on the HLHS patient network, and Sim 8 to the blunted inflow waveform (a) imposed on the HLHS and (b) imposed on DORV.
HLHS Early FALD Late FALD DORV Sim 5 Sim 6 Sim 7 Sim 8a Sim 8b
Systolic Pressure 12.6 12.7 12.9 12.7 14.1 14.1 12.4 11.1 11.0
Diastolic Pressure 7.9 7.9 7.8 7.0 7.0 6.8 7.9 9.8 9.7
Pulse 4.7 4.8 5.1 5.7 7.1 7.3 4.5 1.3 1.3
Average Pressure 10.0 10.1 10.1 9.7 10.6 10.5 10.2 10.5 10.4
@@ 3.1.1 Patient-specific inflows
The pressure and flow of the inlet vessel (vessel 0) differ significantly between the two patients. The HLHS patient has a smoother inflow curve, but the magnitude and frequency of reflective waves increase with the severity of FALD. The incident pressure wave becomes steeper and the increase in wave speed causes the reflected wave to coincide with the systolic peak; as a result, the pressure wave loses the dicrotic notch as FALD progresses.
@@
Figure 6: Pressure and flow predictions for vessels 0 and 13 in the hepatic artery vasculature in a DORV, HLHS, and FALD HLHS patient. Systolic pressures are increased under FALD conditions and in this particular terminal vessel (13) the maximum flow increases under FALD conditions.
A similar analysis is performed for the hepatic artery network. Table 5 shows the predicted pressure and flow in the hepatic artery network. The HLHS patient has a higher systolic and pulse pressure compared to the DORV patient. There is a slight increase in pressures under FALD conditions for the HLHS patient.
We selected two representative vessels (0, 13) in the vasculature of the liver artery to compare flow and pressure predictions between patients (Figure 6). Vessel 0 is the hepatic artery, the network’s root vessel, while vessel 13 is a terminal vessel. Comparing the healthy HLHS and DORV patients, the HLHS has increased systolic, pulse, and average pressures. Systolic pressures are increased as the degree of FALD increases,with a slight decrease in diastolic pressures. This results in an increase of pulse pressures.
Table 5: Predicted pressure and flow in the hepatic artery network of the DORV and HLHS patients. All values are given in mmHg. Sim 5 refers to the HLHS inflow imposed on the DORV network, Sim 6 to the DORV inflow imposed on HLHS network, and Sim 7 to the DORV parameters imposed on the HLHS network.
HLHS Early FALD Late FALD DORV Sim 5 Sim 6 Sim 7
Systolic Pressure 95.3 99.5 99.8 95.1 89.9 104.3 97.3
Diastolic Pressure 83.2 82.9 82.7 83.4 83.2 83.2 83.4
Pulse 12.1 16.6 17.1 11.7 6.7 21.1 13.9
Average Pressure 88.2 88.9 88.8 88.2 86.7 93.8 90.3
3.1.2 Impact of inflow
Imposing the HLHS (Figure S1) and DORV (Figure S2) inflow on the DORV and HLHS networks, was done to study impacts of the inflow waveform shape. Independent of the inflow waveform results show that simulations with the DORV geometry and parameters (i.e., more compliant vessels) lead to smaller pressure build-up, the pressure is lower in the main hepatic artery (Table 5). In the small hepatic artery vessels, the pressures were similar for both patients and inflow. The HLHS inflow is higher, as a result, more flow is pushed through the hepatic arteries, independent of the network. A larger hepatic artery flow results in more flow to the smaller vessels. In the portal network, the higher compliance of the DORV network generates more reflective waves. In this network we observe a significant phase shift in the smaller vessels. Again, similar results were obtained with both inflows.
We also conducted a simulation where the DORV parameterization was imposed on the HLHS geometry to explore impacts of the patient parameterization (Figure S3). Even with the same parameters, similar trends are observed in both flow and pressure predictions (Tables 4 and 5). In the DORV network, the flow maintains a consistent shape and pulsatility, whereas in the HLHS network, the flow becomes flatter in terminal vessels and the pressure are lower in the portal vein network. The main difference is that DORV network has slightly less stiff vessels, as a result wave-propagation speed is slower impacting the arrival time of the reflected waves.
Finally, we imposed a blunted waveform in the portal vein network [chavan] (Figures S4, S5). This waveform shows significant pretense of secondary oscillations resulting from wave reflection generated through the arterial network. Imposing this waveform on the HLHS and DORV network, generates similar flow and pressure amplitudes in both geometries (Table 4). Again, the more compliant DORV vasculature augments secondary flow generating more reflections.
@@ 3.2 Wall Shear Stress Approximation
Graphical representations of WSS over one cardiac cycle at the midpoint of the six previously selected vessels are shown in Figure 7. Inlet vessels (vessel 0) of both vasculatures show the most significant difference in WSS between HLHS and DORV patients. In the hepatic network, it is clear to see that as FALD progresses, WSS values increase as well. The portal network is not as telling, with only small changes, both increases and decreases, as FALD progresses. The DORV patient tends to have higher WSS values in the downstream portal vasculature.
Table 6: Average WSS in the selected DORV and HLHS patient vessels under normal and FALD conditions. Values are given in units of Pa.
@@
Hepatic Artery (0) 9.61 9.60 9.65 6.50
Hepatic Artery (13) 1.89 4.46 4.43 2.96
Portal Vein (0) 0.05 0.06 0.06 0.04
Portal Vein (54) 0.08 0.07 0.07 0.08
Portal Vein (58) 0.08 0.08 0.08 0.10
Portal Vein (69) 0.04 0.04 0.04 0.6